Fermi National Laboratory


Accelerator Update for January 1, 2006 - January 4, 2006

December 21 st, 2005 was the last day of the Accelerator Updates for the year. Most of us went off on vacation or took some time off, but the Accelerators kept running. In fact it was a record week for Tevatron integrated luminosity for the time period from 12/26/05 – 01/02/06 as we integrated about 24,700 inverse nanobarns. We also integrated about 21,000 inverse nanobarns the week prior. Things actually ran fairly smoothly between December 21 st and January 1 st.

We had about 8 stores in the Tevatron with luminosities ranging from 92E30 to 155E30. There were a few quenches at the end of stores, but no major Tevatron downtime. There was some NuMI chiller work at the beginning of this period, but they ran pretty smoothly throughout the entire period. This was true for MiniBooNe and Meson Center (MIPP). Pbar stacked and transferred beam to the Recycler for stashing and later transfers to the Tevatron. There were problems with kickers, bunchers, target rotation, power supplies, chillers and other equipment, but no extended down times were encountered.

Sunday January 1

The midnight shift began with Operations monitoring store #4574, the Recycler Ring (RR) stash at 79.5 E10, Antiproton Source (Pbar) stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the Recycler into shot setup at 6:58 AM to stash antiprotons.

Pbar returns to stacking at 7:47 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 7:39 AM to stash antiprotons.

Pbar returns to stacking at 8:26 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 1:31 PM to stash antiprotons.

Pbar returns to stacking at 2:13 PM, beam also returns to MCenter.

The MI-52 kicker (KPS5S) tripped off at 4:36 PM and won’t reset, experts were called to investigate. This kicker prohibits PBar from stacking.

Monday January 2

The midnight shift began with Operations monitoring store #4574, the Recycler Ring (RR) stash at 268.1 E10, Pbar stack at 42.2 E10, and with MiniBooNE, NuMI, and MCenter taking beam. MI-52 kicker repair is holding off stacking.

The MI-52 kicker was repaired by 12:53 AM, PBar has resumed stacking.

Store 4574 was terminated at 3:06AM.

Shot setup begins at 3:19AM.

Operations established store 4575 at 5:10 AM with an initial luminosity of 164.99 E30.

Operations put the Recycler into shot setup at 4:59 AM to stash antiprotons.

Pbar returns to stacking at 6:08 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 10:03 AM to stash antiprotons.

Pbar returns to stacking at 10:50 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 2:24 PM to stash antiprotons.

Pbar returns to stacking at 3:05 PM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 6:41 PM to stash antiprotons.

Pbar returns to stacking at 7:26 PM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 11:28 PM to stash antiprotons.

Pbar returns to stacking at 12:09 AM, beam also returns to MCenter.

Tuesday January 3

The midnight shift began with Operations monitoring store #4575, the Recycler Ring (RR) stash at 250.5 E10, Pbar stack at 6.7 E10, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the Recycler into shot setup at 4:03 AM to stash antiprotons.

Pbar returns to stacking at 4:52 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 10:08 AM to stash antiprotons.

Pbar returns to stacking at 10:49 AM, beam also returns to MCenter.

Store 4575 was terminated at 1:08 PM.

CDF and D0 went into their collision halls at 1:18 PM

Operations put the Recycler into shot setup at 2:08 PM to stash antiprotons.

At 3:30 PM we set a new record of 350.6 E10 for the recycler stash, the previous record was 318 E10.

CDF and D0 accesses were completed at 3:30 PM.

Pbar returns to stacking at 3:55 PM, beam also returns to MCenter.

At 4:26 PM the Tevatron (Tev) quenched at F4.

At 6:30 PM the Tev was turned off to inspect the F17 kicker. This means no beam anywhere except Linac and Booster.

At 8:00 PM it was determined that F17 kicker #1 was leaking fluid and will have to be replaced.

Wednesday January 4

The midnight shift began with the Recycler Ring (RR) stash at 331.8 E10, Pbar stack at 50 E10. The F17 kicker replacement is in progress. No beam anywhere due to kicker access.

At 3:00 AM, another access found that the F17 collimator has a nice line burned into it. Since the collimator is next to the kicker it will be replaced also.

At 6:30 AM power was lost to the A0 compressor room. A breaker was tripped at the Kautz Road substation. Experts are investigating.

The Plan for Wednesday

Replace the F17 kicker and collimator, do some general maintenance while the beam is off. Try to resume normal Operations sometime this evening.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/4/2006   email Fermilab